Lindley Corp.'s stock price at the end of last year was $22, and its book value per share was $25.00. What was its market/book ratio?

Select the correct answer.

a. 0.81

b. 0.95

c. 0.67

d. 0.74

e. 0.88

Northwest Lumber had a profit margin of 10%, a total assets turnover of 1.5, and an equity multiplier of 1.8. What was the firm's ROE?

Select the correct answer.

a. 26.91%

b. 26.73%

c. 27.00%

d. 26.82%

e. 27.09%

Hutchinson Corporation has zero debt - it is financed only with common equity. Its total assets are $460,000. The new CFO wants to employ enough debt to bring the debt/assets ratio to 40%, using the proceeds from the borrowing to buy back common stock at its book value. How much must the firm borrow to achieve the target debt ratio?

Select the correct answer.

a. $184,014.80

b. $183,985.20

c. $183,955.60

d. $183,970.40

e. $184,000.00

Rappaport Corp.'s sales last year were $355,000, and its net income after taxes was $23,000. What was its profit margin on sales?

Select the correct answer.

a. 6.63 %

b. 6.48 %

c. 6.58 %

d. 6.68 %

e. 6.53 %

Bostian, Inc. has total assets of $560,000. Its total debt outstanding is $185,000. The Board of Directors has directed the CFO to move towards a debt-to-assets ratio of 55%. How much debt must the company add or subtract to achieve the target debt ratio?

Select the correct answer.

a. $123,190

b. $123,063

c. $123,254

d. $123,000

e. $123,127

Heaton Corp. sells on terms that allow customers 45 days to pay for merchandise. Its sales last year were $415,000, and its year-end receivables were $60,000. If its DSO is less than the 45-day credit period, then customers are paying on time. Otherwise, they are paying late. By how much are customers paying early or late? Base your answer on this equation: DSO - Credit period = days early or late, and use a 365-day year when calculating the DSO. A positive answer indicates late payments, while a negative answer indicates early payments.

Select the correct answer.

a. 12.93

b. 11.21

c. 7.77

d. 14.65

e. 9.49
Answer #1

1. Market/book ratio = Stock price / Book value per share

= 22 / 25

= 0.88

correct option: e. 0.88

2. ROE (Return on Equity) = Net Profit / Equity

which can be rewritten as: ROE = (Net Profit / Sales) * (Sales / Total assets) * (Total Assets / Equity)

ROE = Net profit margin * Total asset turnover * Equity Multiplier

ROE = 0.1 * 1.5 * 1.8

ROE = 0.27 = 27%

correct option: c. 27.00%

3. Total assets = $460,000

Required debt/assets ratio = 40%

Required debt to be borrowed = Required debt to asset ratio * Total assets

= 40% * 460000

= $184,000

correct option: e. $184,000.00

4. Profit margin on sales = (Net Income after taxes / Sales) * 100

= (23,000 / 355,000) * 100

= 0.0648 * 100

= 6.48%

correction option: b. 6.48%

5. Total assets = $560,000

Required debt/asset ratio = 55%

Required Debt (as per ratio) = Required debt/asset ratio * Total assets

= 55% * 560000

= $308,000

Additional debt required = Required debt - Current debt outstanding

= 308000 - 185000

= $123,000

correction option: d. $123,000

6. DSO (Days Sales outstanding) = (year-end receivables / Sales) * 365

= (60000 / 415000) * 365

= 52.77 days

Since, DSO > 45 days, means customers are paying late

No. of days late = DSO - 45

= 52.77 - 45

= 7.77

correct option: c. 7.77